A 200-foot tall monument is located in the distance. From a window in a building, a person determines that the angle of elevation to the top of the monument is 14°, and that the angle of depression to the bottom of the tower is 2°. How far is the person from the monument? (Round your answer to three decimal places.)
Answer by math_tutor2020(3835)


Answer: 703.609 feet (approximate)


Explanation

Let's start off with a rough sketch of the diagram

The key points are
A = base of the building
B = location of the person's eye
C = base of the monument
D = location on monument at the same height level as the person's eye
E = top of the monument

We have these variables
x = distance from B to D = distance from A to C
y = distance from C to D

Note how segment DE is 200-y feet long so that CE = CD+DE = y + (200-y) = 200

Let's focus on triangle BCD.
tan(angle) = opposite/adjacent
tan(angle DBC) = CD/BD
tan(2) = y/x
y = x*tan(2)
This will be used later in a substitution step.

Now focus on triangle DEB.
tan(angle) = opposite/adjacent
tan(angle DBE) = DE/BD
tan(14) = (200-y)/x
tan(14) = (200-x*tan(2))/x ........ substitution
x*tan(14) = 200-x*tan(2)
x*tan(14)+x*tan(2) = 200
x*( tan(14)+tan(2) ) = 200
x = 200/( tan(14)+tan(2) )
x = 703.609019511759 approximately
x = 703.609

Make sure your calculator is in degree mode.